Population: 60 Million


Out of Labour force: 10milion


Unemployed People: 10million


Employed people: 25million


a.Calculate the non-participation rate

b. Calculate the labour force

c. Calculate the unemployment rate

d.calculate the labour force participation rate



1
Expert's answer
2021-12-09T15:47:37-0500

Solution:

a.). Non-participation rate = People out of the labor force "\div" Total adult working population x 100

People out of the labor force = 10 million

Total adult working population = 10 + 10 + 25 = 45 million

Non-participation rate = "(\frac{10}{45}) \times 100\% = 22.2\%"



b.). Labor force = Employed people + Unemployed people

Labor force = 25 + 10 = 35 million

 


c.). Unemployment rate = (Unemployed people "\div" Total labor force) "\times" 100

 Unemployment rate = "(\frac{10}{35}) \times 100\% = 28.6\%"

 


d.). Labor force participation rate = (People in the labor force "\div" Total adult working population) "\times" 100

Labor force participation rate = "(\frac{35}{45}) \times 100\% = 77.8\%"